eviant:

A sports stadium has 49,000 seats. Seats sell for $25 in Section A, $20 in Section B, and $15 in Section C. The number of seats in Section A equals the total number of seats in Sections B and C. Suppose the stadium takes in $1,052,000 from each sold-out event. How many seats does each section have?

loveboy:

Where do you find difficulty ?

loveboy:

You there ?

eviant:

I can't find the total amount of seats in each section

mhchen:

Let A = number of seats in A B = number of seats in B C = number of seats in C A = B + C A + B + C = 49,000 25A + 20B + 15C = 1,052,000 Do you know how to solve system of 3 equations?

eviant:

@dude @Vocaloid

Hero:

@eviant, @mhchen helped you out. You have that A = B + C. Substitute B+C in place of A for the 2nd and 3rd equations.

eviant:

@Hero sorry, I can't see @mchen's answer, did he reply directly to my post?

Hero:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @mhchen Let A = number of seats in A B = number of seats in B C = number of seats in C A = B + C A + B + C = 49,000 25A + 20B + 15C = 1,052,000 Do you know how to solve system of 3 equations? \(\color{#0cbb34}{\text{End of Quote}}\)

eviant:

@mhchen @Hero I do not know how to solve them correctly

mhchen:

Have you tried solving a system of 2 equations using substitution? You can use the same method for a system of 3 equations. We have A = B + C We can use that to rewrite the 2nd equation as (B + C) + B + C = 49,000 We can also rewrite the 3rd equation as 25(B+C) = 20B + 15C = 1,052,000 Now you have a system of 2 equations, so you can solve for B and C right?